# WiS Platform: Enhancing Evaluation of LLM-Based Multi-Agent Systems Through Game-Based Analysis

### v1

@@ 6 Conclusion
We develop a novel platform for the game "Who is Spy", providing a versatile environment that facilitates in-depth exploration of model capabilities in attack, defense, reasoning, and deception. Through rigorous experiments and a comprehensive benchmark, our platform has demonstrated its effectiveness in distinguishing multi-agent abilities in complex interactive environments. The dynamic leaderboard, alongside the ease of creating custom agents and utilizing open-source models, offers a significant tool for advancing research in multi-agent system evaluation. Ultimately, we hope our platform will serve as a foundation for further studies into the behaviors and ability of large language models in multi-agent systems.
7 Limitations
Our work introduces a multi-agent game platform for the game "Who is spy". More games will be integrated into our platform in the future. Due to limitations in available API resources, some open-source models have not been tested yet, a broader range of models can be used for the platform’s agents to achieve corresponding results in subsequent developments.
Appendix A Detailed rules

### v2

@@ 7 Conclusion
We develop a novel platform for the game "Who is Spy", providing a versatile environment that facilitates in-depth exploration of model capabilities in attack, defense, reasoning, and deception. Through rigorous experiments, our platform has demonstrated its effectiveness in distinguishing multiagent abilities in complex interactive environments. We hope our platform will serve as a foundation for further studies into the behaviors and ability of large language models in multi-agent systems.
Appendix A Empirical User Research
A.1 Participant Overview
We successfully organized a competition on the Wis platform. A total of 200 participants engaged in the competition, utilizing 204 distinct AI agents across 15,893 game rounds. The competition fostered active discussions and creative strategies among participants.
A.2 Model Selection Trends
The distribution of models used in the competition showed a preference for open-access or lower-cost options. The most commonly used models were DeepSeek (59 agents), GPT-based models (44 agents), QWEN (33 agents), Gemini (2 agents), and Claude (1 agent). The relative scarcity of premium models (e.g., Claude) suggests cost sensitivity among participants.
A.3 Agent Strategy and Model Substitution
An interesting observation emerged from the top performing agents, where discrepancies were observed between the declared and actual models used. Specifically:
  • Some users declared GPT-3.5 but actually employed Doubao, likely to optimize cost efficiency while confusing other users.
  • An agent listed as using Qwen-vl-plus was in fact utilizing glm-plus.
  • Other top-performing agents adhered to their declared models, including Qwen-Max, Claude Sonnet 3.5 , and DeepSeek-V3.
Notably, defensive measures against prompt attacks varied, with some high-ranking agents implementing customized history template formats, keyword filtering, and length restrictions to mitigate injection risks.
A.4 Performance and Game Dynamics
An analysis of top-ranking agents (accessible at https://whoisspy.ai/#/competitionDetail?&id=4&muxsearch=%7B%22type%22%3A%22rank%22%7D) revealed intriguing patterns in scoring. The score distribution was notably tight among participants ranked 2nd to 34th, while the 1st-place agent significantly outperformed others. Further investigation is needed to determine the factors contributing to this score gap, such as strategy optimization, response consistency, or superior prompt engineering.
Regarding win-rate rankings, the top-performing models in terms of victory frequency were:
  • GPT-based models
  • QWEN
  • Claude
  • DeepSeek
This ranking aligns with the broader model usage statistics, suggesting that while cost-effective models dominated in quantity, premium models still achieved competitive win rates.
A.5 Key Findings and Future Directions
  • The high engagement level, despite technical and strategic challenges, indicates strong interest in AI-driven social deduction games.
  • The choice of models was influenced by cost, and free / open source models were more prevalent than premium models.
  • Strategy adaptation played a crucial role, with some users leveraging alternative models or defensive measures against prompt attacks.
  • Score distribution highlights the need for further analysis of key performance differentiators.
  • The adaptation of the strategy played a crucial role, with some users taking advantage of alternative models or defensive measures against prompt attacks.
These insights provide a foundation for refining future competitions, improving accessibility, and further investigating strategic AI behavior in social deduction scenarios.
Appendix B Detailed rules
@@
Appendix C Experimental Setup
All experiments were conducted on our platform, where we deployed well-known open-source models as the default intelligent agents. We evaluated ten publicly available open-source models, each experiment was repeated over 90 times, ensuring consistency in the key generation parameters and sampling algorithms across all models. We conducted an analysis of the capabilities among multiple agents, where each agent utilized the same model, the sole distinction lay in the adoption of our proposed strategy.
In these experiments, we ensured the fairness of the experimental conditions, we ensured that each model played both the roles of an spy agent and a civilian an equal number of times, with each experiment being repeated more than 24 times. We also ensured that in the experiment, each agent utilized identical code, differing only in their base models. In the attacking, defense, and reasoning experiments, all models except for one specialized model used the same code as used in the main experiment.